# ShapeLLM: Universal 3D Object Understanding for Embodied Interaction

### v2

@@ Appendix D Training details
ShapeLLM  All experiments were conducted using 8 $\times$ 80G A800 GPUs with a BF16 data type. During the multimodal alignment stage, we train our model for one epoch with a batch size 256 and a learning rate 2e-3. During the instruction tuning stage, we train our model for one epoch with a batch size of 128 and a learning rate 2e-5. Throughout both stages, we employ flash-attention [26], the AdamW [104] optimizer, and a cosine learning rate scheduler [103]. For the entire training process, the 7B and 13B models require approximately 10 and 20 hours, respectively. Further details regarding hyper-parameters are documented in Tab. 13.
Appendix E Additional Related Work
E.1 3D Representation Learning
@@ Appendix E Additional Related Work
Research on 3D Representation Learning encompasses various methods, including point-based [126, 127], voxel-based [112], and multiview-based approaches [143, 58]. Point-based methods [131, 40] have gained prominence in object classification [164, 151] due to their sparsity yet geometry-informative representation. On the other hand, voxel-based methods [30, 176, 130] offer dense representation and translation invariance, leading to a remarkable performance in object detection [23] and segmentation [175, 3]. The evolution of attention mechanisms [153] has also contributed to the development of effective representations for downstream tasks, as exemplified by the emergence of 3D Transformers [40, 102, 111]. Notably, 3D self-supervised representation learning has garnered significant attention in recent studies. PointContrast [165] utilizes contrastive learning across different views to acquire discriminative 3D scene representations. Innovations such as Point-BERT [180] and Point-MAE [119] introduce masked modeling [60, 31] pretraining into the 3D domain. ACT [36] pioneers cross-modal geometry understanding through 2D or language foundation models such as CLIP [134] or BERT [31]. Following ACT, ReCon [129] further proposes a learning paradigm that unifies generative and contrastive learning. Additionally, leveraging foundation vision-language models like CLIP [36, 134] has spurred the exploration of a new direction in open-world 3D representation learning. This line of work seeks to extend the applicability and adaptability of 3D representations in diverse and open-world/vocabulary scenarios [122, 181, 33, 32, 106].
Appendix F Future Works
@@
ShapeLLM has made significant progress in advancing 3D shape understanding and embodied perception through MLLMs. Future endeavors aim to scale up embodied understanding training using datasets larger than GAPartNet [47], potentially leading to open-vocabulary part-level comprehension, including 6-DoF pose estimation. Excitingly, there is a vision to establish a unified framework capable of comprehending not only 3D shapes but also entire 3D scenes. To enhance real-world applications on robots, a promising approach involves a robotics co-design that effectively connects 3D representations with downstream language-based tasks [72, 77]. Additionally, addressing efficiency for real-time deployment is crucial, emphasizing techniques like model compression [35, 75, 182, 184, 183].

### v3

@@ 0.B.3 ChatGPT Costs
Appendix 0.C Implementation details
ReCon++  Following the standard ViT [39] architecture, we design four different model structures consistent with prior work [135, 97, 203]. The model parameters are shown in Tab. 14. Following OpenShape [97], we employ four datasets as pretraining data, namely Objaverse [30], ShapeNet [13], ABO [23], and 3D-FUTURE [44]. Each point cloud sample has a size of 10,000$\times$6, where the first three dimensions represent $xyz$ coordinates, and the latter three dimensions represent $rgb$ values.
@@ Appendix 0.D Training details
ShapeLLM  All experiments were conducted using 8 $\times$ 80G A800 GPUs with a BF16 data type. During the multimodal alignment stage, we train our model for one epoch with a batch size 256 and a learning rate 2e-3. During the instruction tuning stage, we train our model for one epoch with a batch size of 128 and a learning rate 2e-5. Throughout both stages, we employ flash-attention [27], the AdamW [107] optimizer, and a cosine learning rate scheduler [106]. For the entire training process, the 7B and 13B models require approximately 10 and 20 hours, respectively. Further hyper-parameters are documented in Tab. 16.
Appendix 0.E Additional Related Work
0.E.1 3D Representation Learning
@@ Appendix 0.E Additional Related Work
Various methods have been proposed to tackle 3D Representation Learning, including point-based [132, 133], voxel-based [115], and multiview-based approaches [149, 61]. Point-based methods [137, 41] have gained prominence in object classification [170, 157] due to their sparsity yet geometry-informative representation. On the other hand, voxel-based methods [31, 183, 136] offer dense representation and translation invariance, leading to a remarkable performance in object detection [24] and segmentation [182, 3]. The evolution of attention mechanisms [159] has also contributed to the development of effective representations for downstream tasks, as exemplified by the emergence of 3D Transformers [41, 105, 114]. Notably, 3D self-supervised representation learning has garnered significant attention in recent studies. PointContrast [171] utilizes contrastive learning across different views to acquire discriminative 3D scene representations. Innovations such as Point-BERT [186] and Point-MAE [124] introduce masked modeling [63, 32] pretraining into the 3D domain. ACT [37] pioneers cross-modal geometry understanding through 2D or language foundation models such as CLIP [138] or BERT [32]. Following ACT, ReCon [135] further proposes a learning paradigm that unifies generative and contrastive learning, which can be applied to both single-modal or cross-modal settings. Additionally, leveraging foundation vision-language models like CLIP [37, 138] has spurred the exploration of a new direction in open-world 3D representation learning. This line of work seeks to extend the applicability and adaptability of 3D representations in diverse and open-world/vocabulary scenarios by distilling the open-world knowledge within foundation models [127, 188, 33, 34, 109, 42], with which it is now possible to perceive the 3D physical scenes using human languages.
Appendix 0.F Future Works
@@
ShapeLLM has made significant progress in advancing 3D shape understanding and embodied perception through MLLMs. Future endeavors aim to scale up embodied understanding training using datasets larger than GAPartNet, potentially leading to open-vocabulary part-level comprehension, including 6-DoF pose estimation. To this end, the first possibility is to empower the training data and benchmarking data with more advanced MLLMs such as GPT4-o [121], which are more human-aligned intelligent agents [169, 128]. Excitingly, there is a vision to establish a unified framework capable of comprehending not only 3D shapes but also entire 3D scenes. To enhance real-world applications on robots, a promising approach involves a robotics co-design that effectively connects 3D representations with downstream language-based tasks [74, 80, 20]. Additionally, addressing efficiency for real-time deployment is crucial, emphasizing techniques like model compression [38, 78, 189, 191, 190].
Figure 15: Additional Visualization example of 3D MM-Vet Q&A pairs.
Figure 16: Additional Visualization example of 3D MM-Vet Q&A pairs.